OpenStudy (anonymous):

The perimeter of a rectangle is 42ft. The length is 5ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. 2x + 2y = x=y + ? what is the length ? ft

OpenStudy (anonymous):

2x+2y is the perimeter. 2x+2y=42 the length is 5 greater than the width so x=y+5 We solve by multiplying the second equation by -2 and adding the two equations. 2x+2y=42 -2x+2y=-10 4y=32 y=8 x=13